OpenStudy (anonymous):

If it requires 38.5 milliliters of 0.85 molar nitric acid to neutralize 20.0 milliliters of barium hydroxide, solve for the molarity of barium hydroxide. Show all of the work used to solve this problem. Unbalanced equation: Ba(OH)2 + HNO3 Ba(NO3)2 + H2O ( I already balanced the equation)

OpenStudy (jfraser):

the reaction's not balanced

OpenStudy (anonymous):

Ba(OH)2 + 2 HNO3 = Ba(NO3)2 + 2 H2O

OpenStudy (anonymous):

we are given Ba(OH)2: Vol = 0.02 (have to write in litres), M (conc.) = y (what we're finding out) HNO3: Vol. = 0.0385, M = 0.85 n (moles) = MxV n(HNO3) = 0.85 x 0.0385 = 0.032725 The mole ratio is 2 : 1 since HNO3 has one H+ and Ba(OH)2 has two OH- Therefore n(Ba(OH)2) = 0.032725 / 2 = 0.0163625 M = n / v therefore y = 0.0163625 / 0.020 = 0.818125 roughly = 0.82M :) any doubt??

OpenStudy (anonymous):

Thankk you so much

OpenStudy (anonymous):

yw but its a humbly request frnd plz dont post qn on math sec as i really dont belong to this subject so it just screw wen i see this sec :P

OpenStudy (anonymous):

OHhh I see sorry! I still thought I was in chemistry. it must have took me here